# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v2

@@ 5 Examples of 8-quadric configurations
We start with a sphere as quadric $\mathbf{{S}}_{0}$. Then we choose 3 planes $p_{0}^{i}$ intersecting the sphere in 3 rings $\mathbf{{R}}_{0}^{i}$ such that each pair of rings meet in a pair of points. The corresponding cones touching the sphere along those rings are taken as quadrics $\mathbf{{T}}^{i}$. Now for each pair the intersection $\mathbf{{T}}^{i}\cap\mathbf{{T}}^{j}$ consists of two conics, see figure 7. We will choose one of them as a degenerate quadric $\mathbf{{S}}_{k},\{i,j,k\}=\{1,2,3\}$. This situation can be pictured as a limit of thin ellipsoids in ring contact with $\mathbf{{T}}^{i},\mathbf{{T}}^{j}$ whose planes of ring contact $p_{k}^{i},p_{k}^{j}$ finally converge to a single plane, which contains the chosen conic of the intersection.
@@
There are now two possibilities for such a triple of conics $\{\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}\}$, one leading to a regular $T_{0}$ and one not: Either the 3 conics meet in a point pair, which lies in the intersection $\mathbf{{T}}^{1}\cap\mathbf{{T}}^{2}\cap\mathbf{{T}}^{3}$ of all the three cones, or they form a conical octahedron.
In the latter case the final quadric $\mathbf{{T}}^{0}$ is regular (directly given by lemma 3.4). Since each quadric $\mathbf{{S}}_{i}$ degenerates to a conic, all 3 rings $\mathbf{{R}}_{i}^{k}$ on it also collapse to this conic and each axis $\Lambda_{ij}^{0k}$ is the intersection of two ring planes. In the former case, on the other hand, the final quadric $\mathbf{{T}}^{0}$ is a point pair. This case, however, cannot as easily be subsumed into the logic of section 3.3 since $\Lambda^{01}_{23}=\Lambda^{02}_{13}=\Lambda^{03}_{12}$. To picture this in terms of a limit, we consider a sequence of shrinking ellipsoids, the $\mathbf{{S}}_{i}$ getting flatter and $\mathbf{{T}}^{0}$ getting thinner and finally collapsing to a line segment connecting the two points, all planes through either point ultimately being tangent planes enveloping the ellipsoid in the limit.
Although these configurations are quite elementary in 3D their slices with a plane yield nontrivial theorems in 2D. In particular consider the plane $p$ through the ring pointsof the 3 cones: It intersects the cones in 6 lines tangential to a circle being the intersection of the sphere $\mathbf{{S}}_{0}$ with this plane. In the case of three concurrent conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ they intersect the plane $p$ in 6 points which are joined to a hexagon by the tangent lines. This leads to a Brianchon configuration. In the case of three non-concurrent conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ they intersect the plane $p$ in 6 points which split up into two triangles circumscribed to the circle of intersection. This leads to a Poncelet configuration. A slice with a general plane gives the double contact theorem and the 4-conic theorem from [EMCT74] . Note that a single 3D-configuration contains Brianchon figures in 4 different ways such that the diagonals form a complete 4-point.
Figure 8: Left: Three concurrent conics of the conical cube induce a hexagon circumscribed to a circle in the plane $p$ through the ring pointsof the cones and thus proves the Brianchon theorem. Right: In the case of three non-concurrent conics the 6 points split into two triangles. This figure proves the Poncelet theorem.
@@
In the next example we consider three spheres $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ such that no sphere is completely contained in another one. All spheres intersect the plane at infinity in the same imaginary circle $\mathbf{{S}}_{0}$, the absolute of Euclidean geometry. Thus each sphere $\mathbf{{T}}^{i}$ is in ring contact with $\mathbf{{S}}_{0}$, where the quadric $\mathbf{{S}}_{0}$ as an imaginary degenerated conic represents at the same time the rings $\mathbf{{R}}_{0}^{i}$. We could proceed as in the previous example by choosing as quadrics $\mathbf{{S}}_{k}$ the intersection of the spheres $\mathbf{{T}}^{i}\cap\mathbf{{S}}_{0}$, which are circles (besides the universal imaginary circle at infinity) and by slicing with a plane through the centers of the 3 spheres we would obtain again a Brianchon-like theorem, which turns out to be the theorem of the radical center of 3 circles. Instead we choose cones enveloping each pair of spheres. If the spheres are disjoint there are 6 such cones in total (dual to the 6 conics in example 5.1) and again there are two cases for the choice of a triplet of cones. In one case the cones share 2 tangent planes which are also common tangent planes of the 3 spheres. Thus the final quadric $\mathbf{{T}}^{0}$ is this pair of planes and the rings $\mathbf{{R}}_{i}^{0}$ are 3 line pairs. This 3D-configuration was already used by G. Monge (according to [DW92]p.154) in order to prove a theorem nowadays bearing his name: The three pairs of external tangents of three circles meet in three points on a line.
In the other case the cones do not admit common tangent planes. Now the rings $\mathbf{{R}}_{i}^{0}$ are obtained by intersecting each cone $\mathbf{{S}}_{i}$ with the corresponding plane $p_{i}^{0}$ which is the join of a pair of vertical axis. The final quadric is again guarantied by lemma 3.4. A plane slice through the centers of the spheres shows that 3 pairs of internal tangents of 3 circles are circumscribed to a conic. This theorem seems to be known only for the case, where the final conic degenerates into a point pair, see [Bak36],[Sal17].
@@
Consider 3 planes $p_{0}^{i}$ tangent to a quadric $\mathbf{{S}}_{0}$ at points $P_{0}^{i}$. In order to see the rings $\mathbf{{R}}_{0}^{i}$ as real objects we choose here a regulus as quadric $\mathbf{{S}}_{0}$, thus each ring is a line pair contained in the regulus. This is the case referred to as X-contact in section 2.2, where the ring pointis incident to the ring plane. We want to picture these line pairs as ‘quadrics’ $\mathbf{{T}}^{i}$, though such quadrics don’t actually exist. They are in fact double planes when considered point-wise and double points when considered plane-wise, but for the sake of this discussion we shall identify the $\mathbf{{T}}^{i}$ with the line pairs (X’s of contact). We can imagine them as limits of flat hyperboloids touching the regulus along the line pair of intersection $\mathbf{{S}}_{0}\cap p_{0}^{i}$ with a tangent plane.
@@
Each couple of X’s of contact $\mathbf{{T}}^{i},\mathbf{{T}}^{j}$ meet in a pair of points and are contained in a pair of planes since $\mathbf{{T}}^{i},\mathbf{{T}}^{j}\subset\mathbf{{S}}_{0}$. We choose the plane pairs to be the quadrics $\mathbf{{S}}_{k}$. Then the final quadric is the double plane spanned by the three points $P_{0}^{1},P_{0}^{2},P_{0}^{3}$ we started with and the rings $\mathbf{{R}}_{i}^{0}$ are the 3 sides of this triangle, each side considered as a double line, see figure 10.
@@
The configuration described above is well known: The 3 plane pairs $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ form a hexahedron which is inscribed to a regulus such that 6 of the edges (a skew hexagon) lie in the regulus. Thus a plane slice of this configuration yields a hexagon on a conic and it shows that opposite sides of such a hexagon (corresponding to the plane pairs $\mathbf{{S}}_{i}$) meet in 3 points which lie on a line (corresponding the the plane $\mathbf{{T}}^{0}$).
@@
Figure 10: The proof of the Pascal theorem by a hexahedron inscribed to a regulus was found by G. Dandelin [Dan26]. The Pascal axis is the intersection of the planar slice (bottom) with the ideal plane of the hexahedron (top.)
@@
Finally we will look at the same configuration but construct it in reverse order, starting with a double plane $\mathbf{{S}}_{0}$ followed by three plane pairs $\mathbf{{T}}^{i}$, each plane pair meeting in a line contained in $\mathbf{{S}}_{0}$. Each pair of plane pairs intersects in two line pairs. In this special case the conical cube of the 6 line pairs are the opposite edges of a usual projective cube. As in the examples before a triplet of chosen line pairs might be concurrent or not. In the first case the 3 line pairs obviously split into two triples of lines, belonging to oppisite vertices in the cube, see figure 11 left. A planar slice yields Desargues’ theorem. For the other choise the 3 line pairs form a closed zigzag and the planar slice yields Pascal’s theorem. In both cases the axis in the theorem is the intersection with the plane $\mathbf{{S}}_{0}$.
@@
Figure 11: If three line pairs are chosen, which meet in two points (left) their plane slice splits into two perspective triangles and yields Desargues (middle). Otherwise the three line pairs form a zigzag and induce an hexagon in a planar slice and thus Pascal (right).
@@
The above examples show the following common structure: The three quadrics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ give rise to 6 degenerate quadrics either by pairwise intersections (examples 5.1 and 5.3) or by pairwise envelops (example 5.2). Depending on the choice of such a triplet $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ out of these six the final quadric $\mathbf{{T}}^{0}$ is either regular or a pair of points (or planes). This is summarized in figure 12.
@@ 6 Related research and outlook for future work
The research leading to this article has also produced two companion articles which provide a different perspective on the material. [CGN] presents a proof of the 8-conic theorem working within the 5D parameters space of conics. [ACG+] does the same for the 8-quadric theorem within the 9D space of quadrics. The proofs have a similar structure to each other, but the advantages of the 3D parameter-space approach over the 2D approach becomes clear in the final stages of the proofs.
@@
Regarding cube configurations with degenerate conics or quadrics, further conditions are needed to ensure the validity of the main theorems, presented in this article. This can be investigated in parameter space by looking at limits of regular configurations. It seems very worthwhile to us to further develop these techniques making use of the geometric structures of parameter space, as done in [CGN] and [ACG+].

### v3

@@ 5 Examples of 8-quadric configurations
Consult Fig. 7. We start with a sphere (green) as quadric $\mathbf{{S}}_{0}$. Then we choose 3 planes $p_{0}^{i}$ intersecting the sphere in 3 (blue) rings $\mathbf{{R}}_{0}^{i}$ such that each pair of rings meet in a pair of points (white). They form a conical octahedron, featured above in Lemma 3.4. The corresponding cones touching the sphere along those rings are taken as quadrics $\mathbf{{T}}^{i}$ (yellow). Now for each pair of cones the intersection $\mathbf{{T}}^{i}\cap\mathbf{{T}}^{j}$ consists of two conics (red). These six conics meet three at a time in the 8 yellow points, the common intersection $\mathbf{{T}}^{1}\cap\mathbf{{T}}^{2}\cap\mathbf{{T}}^{3}$ of all the three cones. We call this configuration a conical cube, whereby the yellow points represent the vertices and each red conic represents a pair of opposite edges of the cube.
We will choose one conic from each of the three pairs as a degenerate quadric $\mathbf{{S}}_{k},\{i,j,k\}=\{1,2,3\}$. This situation can be pictured as a limit of thin ellipsoids in ring contact with $\mathbf{{T}}^{i},\mathbf{{T}}^{j}$ whose planes of ring contact $p_{k}^{i},p_{k}^{j}$ finally converge to a single plane, which contains the chosen conic of the intersection.
@@
There are now two possibilities for such a triple of conics $\{\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}\}$, one leading to a regular $T_{0}$ and one not: Either the 3 conics meet in a point pair, or they form a conical octahedron. In Figure 7, left, the first possibility is shown by the three red conics that meet in the yellow point in the middle surrounded by the three blue ring conics (with a second meeting point on the opposite side of the sphere), while the second possibility is shown by the remaining three red conics, that surround the point in the middle and extend farther out along the 3 yellow cones and meet pairwise in two points.
In the latter case the final quadric $\mathbf{{T}}^{0}$ is regular (directly given by Lemma 3.4). Since each quadric $\mathbf{{S}}_{i}$ degenerates to a conic, all 3 rings $\mathbf{{R}}_{i}^{k}$ on it also collapse to this conic and each axis $\Lambda_{ij}^{0k}$ is the intersection of two ring planes. In the former case, on the other hand, the final quadric $\mathbf{{T}}^{0}$ is a point pair. This case, however, cannot as easily be subsumed into the logic of section 3.3 since the final 3 axes are identical: $\Lambda^{01}_{23}=\Lambda^{02}_{13}=\Lambda^{03}_{12}$. To picture this in terms of a limit, we consider a sequence of shrinking ellipsoids, the $\mathbf{{S}}_{i}$ getting flatter and $\mathbf{{T}}^{0}$ getting thinner and finally collapsing to a line segment connecting the two points, all planes through either point ultimately being tangent planes enveloping the ellipsoid in the limit.
Although these configurations are quite elementary in 3D their slices with a plane yield nontrivial theorems in 2D. In particular, consider the plane $p$ through the ring points of the 3 cones: It intersects the cones in 6 lines tangential to a circle being the intersection of the sphere $\mathbf{{S}}_{0}$ with this plane. In the case of three concurrent conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$, they intersect the plane $p$ in 6 points which form a hexagon circumscribing the circle of intersection with the sphere. This leads to a Brianchon configuration. In the case of three non-concurrent conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ they intersect the plane $p$ in 6 points which split up into two triangles circumscribing the circle of intersection. This leads to a Poncelet configuration. A slice with a general plane gives the double contact theorem and the 4-conic theorem from [EMCT74]. Note that a single 3D-configuration contains Brianchon figures in 4 different ways such that the Brianchon diagonals form a complete 4-point.
Figure 8: Left: Three concurrent conics of the conical cube induce a hexagon circumscribing a circle in the plane $p$ through the ring points of the cones and thus proves the Brianchon theorem. Right: In the case of three non-concurrent conics the 6 points split into two triangles. This figure proves the Poncelet theorem.
@@
In the next example we consider three spheres $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ such that no sphere is completely contained in another one. All spheres intersect the plane at infinity in the same imaginary circle $\mathbf{{S}}_{0}$, the absolute of Euclidean geometry. Thus each sphere $\mathbf{{T}}^{i}$ is in ring contact with $\mathbf{{S}}_{0}$, where the quadric $\mathbf{{S}}_{0}$ as an imaginary degenerate conic represents at the same time the rings $\mathbf{{R}}_{0}^{i}$. We could proceed as in the previous example by choosing as quadrics $\mathbf{{S}}_{k}$ the intersection of the spheres $\mathbf{{T}}^{i}\cap\mathbf{{S}}_{0}$, which are circles (besides the absolute imaginary circle at infinity) and by slicing with a plane through the centers of the 3 spheres we would obtain again a Brianchon-like theorem, which turns out to be the theorem of the radical center of 3 circles.
Instead we choose cones enveloping each pair of spheres. If the spheres are disjoint there are 6 such cones in total (dual to the 6 conics in Example 5.1) and again there are two cases for the choice of a triplet of cones. In one case the cones share 2 tangent planes which are also common tangent planes of the 3 spheres. See Figure 9, left. Thus the final quadric $\mathbf{{T}}^{0}$ is this pair of planes and the rings $\mathbf{{R}}_{i}^{0}$ are 3 line pairs. This 3D-configuration was already used by G. Monge (according to [DW92]p.154) in order to prove a theorem nowadays bearing his name: The three pairs of external tangents of three circles meet in three points on a line.
In the other case the cones do not admit common tangent planes. See Figure 9, right. Now the rings $\mathbf{{R}}_{i}^{0}$ are obtained by intersecting each cone $\mathbf{{S}}_{i}$ with the corresponding plane $p_{i}^{0}$ which is the join of a pair of vertical axes. The final quadric is again guarantied by lemma 3.4. A planar slice through the centers of the spheres shows that 3 pairs of internal tangents of 3 circles are circumscribed to a conic. This theorem seems to be known only for the case, where the final conic degenerates into a point pair, see [Bak36],[Sal17].
@@
Consider 3 planes $p_{0}^{i}$ tangent to a quadric $\mathbf{{S}}_{0}$ at points $P_{0}^{i}$. In order to see the rings $\mathbf{{R}}_{0}^{i}$ as real objects we choose here a regulus as quadric $\mathbf{{S}}_{0}$, thus each ring is a line pair contained in the regulus. Consult Fig. 10. This is the case referred to as X-contact in section 2.2, whereby the ring point is incident to the ring plane. We want to picture these line pairs as ‘quadrics’ $\mathbf{{T}}^{i}$, although the situation is more complex. They are in fact double planes when considered point-wise and double points when considered plane-wise, but in an X-contact pencil we can identify them with the line pairs themselves. We can imagine these ‘quadrics’ as the limit of a sequence of reguli all containing the line pair $\mathbf{{S}}_{0}\cap p_{0}^{i}$, as they get flatter and flatter.
@@
Each pair of line pairs $\mathbf{{T}}^{i},\mathbf{{T}}^{j}$ meet in a pair of points on $\mathbf{{S}}_{0}$ and are contained in a pair of tangent planes since $\mathbf{{T}}^{i},\mathbf{{T}}^{j}\subset\mathbf{{S}}_{0}$. We choose the plane pairs to be the quadrics $\mathbf{{S}}_{k}$. Then the eighth quadric is the double plane spanned by the three points $P_{0}^{1},P_{0}^{2},P_{0}^{3}$ we started with, and the rings $\mathbf{{R}}_{i}^{0}$ are the 3 sides of this triangle, each side considered as a double line, see Figure 10.
@@
The configuration described above is well known. The 3 plane pairs $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ form a hexahedron which is inscribed around a regulus such that 6 of the edges (a skew hexagon) lie in the regulus. Thus a planar slice of this configuration yields a hexagon inscribed in a conic such that opposite sides of the hexagon (corresponding to the plane pairs $\mathbf{{S}}_{i}$) meet in 3 points which lie on a line (the intersection of the slicing plane with the plane $\mathbf{{T}}^{0}$), the Pascal line of the hexagon. This 3D proof of Pascal’s Theorem was already found in 1826 by G. Dandelin [Dan26].
@@
Figure 10: The proof of Pascal’s Theorem using a hexahedron inscribed around a regulus. The Pascal axis is the intersection of the planar slice (bottom) with the ideal plane of the hexahedron (top).
@@
Finally we will look at the same configuration but construct it in reverse order, starting with a double plane $\mathbf{{S}}_{0}$ followed by three plane pairs $\mathbf{{T}}^{i}$, each plane pair meeting in a line contained in $\mathbf{{S}}_{0}$. Each pair of plane pairs intersects in two line pairs. In this special case the conical cube of the 6 line pairs are the opposite edges of a usual projective cube. As in the examples before a triplet of chosen line pairs might be concurrent or not. In the first case the 3 line pairs obviously split into two triples of lines, belonging to opposite vertices in the cube, see Figure 11 left. A planar slice yields Desargues’ theorem. For the other choice, the 3 line pairs form a closed zigzag and the planar slice yields Pascal’s theorem. In both cases the axis in the theorem is the intersection with the plane $\mathbf{{S}}_{0}$.
@@
Figure 11: If three pairs of opposite edges on a hexahedron (a projective cube) are chosen, such that one line from each pair meet in two opposite vertices (left), a planar slice splits into two perspective triangles, proving Desargues’ Theorem (middle). Otherwise the three edge pairs form a zigzag on the cube and yield a hexagon in a planar slice, proving Pascal’s Theorem (right).
@@
The above examples show the following common structure: The three quadrics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ give rise to 6 degenerate quadrics either by pairwise intersections (Examples 5.1 and 5.3) or by pairwise envelopes (Example 5.2). Depending on the choice of such a triplet $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$, one from each pair, the final quadric $\mathbf{{T}}^{0}$ is either regular or a pair of points (or planes). This is summarized in Figure 12.
@@ 6 Related research and outlook for future work
The research leading to this article has also produced two companion articles which provide a different perspective on the material. [CGN] presents a proof of the 8-conic theorem working within the 5D parameters space of conics. [ACG+] does the same for the 8-quadric theorem within the 9D space of quadrics. The proofs have a similar structure to each other, but advantages of the 3D parameter-space approach over the 2D approach become clear in the final stages of the proofs.
@@
Regarding cube configurations with degenerate conics or quadrics, further conditions are needed to ensure the validity of the main theorems, presented in this article. This can be investigated in parameter space by looking at limits of regular configurations. It seems very worthwhile to us to further develop these techniques making use of the projective geometric structures of parameter space, as done in [CGN] and [ACG+].